# Dimension-independent convergence rates of randomized nets using median-of-means

### v2

@@ 5 Tractability under more general randomization
$\lim_{s\to\infty}\sum_{j=1}^{s}\gamma_{j}^{2}2^{\tilde{t}_{j}(2(\alpha+\lambda)+1)}<\infty.$
While increasing $d$ may appear to enhance tractability, parameters $\tilde{t}_{u}$ and $\tilde{t}_{j}$ typically scale with $d$, as in the interlacing construction [4], and unbounded increases in $d$ ultimately violate the tractability condition above. Thus, prudent selection of $d$ is critical for practical implementations.
@@
Suppose $\|f\|_{\alpha,\lambda}<\infty$ for $\alpha\in\mathbb{N}_{0}$ and $\lambda\in(0,1]$. Given a randomization scheme satisfying Assumption 1 with $d<\alpha+\lambda$, we further assume $\gamma_{u}\leqslant\prod_{j\in u}\gamma_{j}$ for $\{\gamma_{j},j\in 1{:}s\}$, $\tilde{t}_{u}\leqslant\sum_{j\in u}\tilde{t}_{j}$ for $\{\tilde{t}_{j},j\in 1{:}s\}$ and
$\lim_{s\to\infty}\sum_{j=1}^{s}\Big{(}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{j}2^{\tilde{t}_{j}}\Big{)}^{\phi}<\infty\quad\text{for}\quad\phi=\frac{\alpha+\lambda+1/2}{\alpha+\lambda-d+1/2}.$ (43)
@@
$\Pr\Big{(}|\hat{\mu}^{(r)}_{\infty}-\mu|^{2}>C_{\eta}2^{-(2\alpha+2\lambda+1-2\eta)m}\Big{)}\leqslant 2^{-r}.$
@@
$\Gamma(m)=1+\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\Big{(}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}2^{\tilde{t}_{u}}\Big{)}^{\phi}\Big{(}(A_{\alpha,\lambda}m+B_{\alpha,\lambda})^{|u|}+2^{\frac{(1-\theta^{\prime})d}{\alpha+\lambda-d}m}D^{|u|}_{\alpha,\lambda,d,\theta^{\prime}}\Big{)}$
@@
$\leqslant$ $\frac{\delta^{\prime}}{\Gamma(m)}\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\Big{(}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}2^{\tilde{t}_{u}}\Big{)}^{\phi}\Big{(}(A_{\alpha,\lambda}m+B_{\alpha,\lambda})^{|u|}+2^{\frac{(1-\theta^{\prime})d}{\alpha+\lambda-d}m}D^{|u|}_{\alpha,\lambda,d,\theta^{\prime}}\Big{)},$
@@
which is bounded as $s\to\infty$ because of equation (43). Our conclusion then follows from Lemma 2 with $\delta=1/16$ and equation (5). ∎
@@
$\sum_{j=1}^{s}\Big{(}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{j}2^{\tilde{t}_{j}}\Big{)}^{\phi}$
$\leqslant$ $\Big{(}\sum_{j=1}^{s}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{j}\Big{)}^{1-\frac{\phi}{1+2(\alpha+\lambda)}}\Big{(}\sum_{j=1}^{s}\gamma^{\frac{1+2d}{\alpha+\lambda+1/2}}_{j}2^{\tilde{t}_{j}(2(\alpha+\lambda)+1)}\Big{)}^{\frac{\phi}{1+2(\alpha+\lambda)}}.$
@@
For the random linear scrambling using base-2 Niederreiter sequences, substituting the bound (42) into equation (43) yields the simplified tractability condition:
$\lim_{s\to\infty}\sum_{j=1}^{s}\Big{(}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{j}j\log_{2}(j)\Big{)}^{\phi}<\infty.$
Since $\phi>1$, this condition is strictly weaker than the requirements in [9, Corollary 3.11] for such digital nets to achieve near $O(2^{-(\alpha+\lambda+1/2)m})$ convergence rates.
In general, when $\gamma_{j}=O(j^{-p})$ and $2^{\tilde{t}_{j}}=O(j^{q})$ for $p\geqslant 0$ and $q\geqslant 0$, $\mu^{(r)}_{\infty}$ attains dimension-independent convergence if
@@ 6 Numerical experiments
In this section, we test our theory across integrands with varying dimensions and decay rates on $\gamma_{u}$. The test integrands have the common form
@@
For each integrand $f_{s,\gamma,\alpha}$, we evaluate the performance of the median estimator $\hat{\mu}^{(r)}_{E}$ for $m=1,2,\dots,16$ under two randomization schemes: the completely random design (CRD) and the random linear scrambling (RLS). The estimator $\hat{\mu}^{(r)}_{E}$ is computed as the sample median of $12r-$ 1 independent realizations of $\hat{\mu}_{E}$, where $\hat{\mu}_{E}$ is defined as in equation (2) with precision $E=64$. The observed integration errors in our tested range of $m$ is much larger than $2^{-E}$, ensuring that $\hat{\mu}^{(r)}_{E}$ closely approximates $\hat{\mu}^{(r)}_{\infty}$. Consequently, the error bounds in Theorem 1, 2 and 3 apply to $\hat{\mu}^{(r)}_{E}$ with probability at least $1-2^{-r}$. We set $r=10$, limiting the failure probability per data point to below $0.1\%$. As a benchmark, we compute the standard RQMC estimator (STD), defined as the average of the $12r-$ 1 $\hat{\mu}_{E}$ realizations used by RLS. The generating matrices $\mathcal{C}_{j}$ used by RLS and STD come from [12].
@@
Finally, we compare the integration errors under RLS and STD. Across the vast majority of test cases, RLS demonstrates superior performance over STD, reinforcing that the median RQMC approach (via RLS) is a more robust alternative to traditional RQMC (via STD). We also note that STD exhibits convergence rates comparable to those of RLS and substantially surpasses the theoretical $O(2^{-(3/2)m})$ rate derived from the root mean squared error when $\gamma=3$ or $4$. This discrepancy arises because each estimator $\hat{\mu}_{\infty}$ in STD is a special case of $\hat{\mu}^{(r)}_{\infty}$ with $r=1$, which satisfies the rapidly converging probabilistic error bounds in Theorems 2 and 3. However, averaging multiple replicates in STD compromises these bounds, as even a single outlier among the replicates can dominate the error and inflate its magnitude. This is the key mechanism that motivates our shift toward median RQMC approaches.
@@ 7 Discussion
A natural extension of our analysis is to integrands with discontinuities in their derivatives or function values. As discussed in Subsection 2.4, the requirement $f\in C^{(\alpha,\dots,\alpha)}([0,1]^{s})$ can be relaxed provided that the weak derivatives satisfy equation (6). Notably, the fractional Vitali variation framework inherently accommodates discontinuities. By developing a theory for non-smooth integrands, our results could be extended to real-world applications in finance and related fields. For instance, [10] showcases an infinite-dimensional integration problem where all ANOVA components are smooth despite discontinuities in the integrand’s derivatives, suggesting that non-smoothness may not fundamentally compromise convergence rates in high-dimensional settings.

### v3

@@ 5 Tractability under more general randomization
$\sum_{j=1}^{\infty}\gamma_{j}^{2}2^{\tilde{t}_{j}(2(\alpha+\lambda)+1)}<\infty.$
While increasing $d$ may appear to enhance tractability, parameters $\tilde{t}_{u}$ and $\tilde{t}_{j}$ typically scale with $d$, as in the interlacing construction [dick:2011], and unbounded increases in $d$ ultimately violate the tractability condition above. Thus, prudent selection of $d$ is critical for practical implementations.
@@
Suppose $\|f\|_{\alpha,\lambda}<\infty$ for $\alpha\in\mathbb{N}_{0}$ and $\lambda\in(0,1]$. Given a randomization scheme satisfying Assumption 1 with $d<\alpha+\lambda$, we further assume $\gamma_{u}\leqslant\prod_{j\in u}\gamma_{j}$ for $\{\gamma_{j},j\in\mathbb{N}\}$, $\tilde{t}_{u}\leqslant\sum_{j\in u}\tilde{t}_{j}$ for $\{\tilde{t}_{j},j\in\mathbb{N}\}$ and
$\sum_{j=1}^{\infty}\Big(\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{j}2^{\tilde{t}_{j}}\Big)^{\phi}<\infty\quad\text{for}\quad\phi=\frac{\alpha+\lambda+1/2}{\alpha+\lambda-d+1/2}.$ (31)
@@
$\Pr\Big(|\hat{\mu}^{(r)}_{\infty}-\mu|^{2}>C_{\eta}\|f\|^{2}_{\alpha,\lambda}2^{-(2\alpha+2\lambda+1-2\eta)m}\Big)\leqslant 2^{-r}.$
@@
[…]
@@
[…]
@@
which is bounded as $s\to\infty$ because of equation (31). Our conclusion then follows from Lemma 2 with $\delta=1/16$. ∎
@@
$\sum_{j=1}^{\infty}\Big(\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{j}2^{\tilde{t}_{j}}\Big)^{\phi}$
$\leqslant$ $\Big(\sum_{j=1}^{\infty}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{j}\Big)^{1-\frac{\phi}{1+2(\alpha+\lambda)}}\Big(\sum_{j=1}^{\infty}\gamma^{\frac{1+2d}{\alpha+\lambda+1/2}}_{j}2^{\tilde{t}_{j}(2(\alpha+\lambda)+1)}\Big)^{\frac{\phi}{1+2(\alpha+\lambda)}}.$
@@
For base-2 Niederreiter sequences, substituting the bound (30) into equation (31) yields the simplified tractability condition:
$\sum_{j=1}^{\infty}\Big(\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{j}j\log_{2}(j)\Big)^{\phi}<\infty.$
Since $\phi>1$, this condition is strictly weaker than the requirements in [Goda2024, Corollary 3.11] for such digital nets to achieve near $O(2^{-(\alpha+\lambda+1/2)m})$ convergence rates under the random linear scrambling.
In general, when $\gamma_{j}=O(j^{-p})$ and $2^{\tilde{t}_{j}}=O(j^{q})$ for $p\geqslant 0$ and $q\geqslant 0$, $\hat{\mu}^{(r)}_{\infty}$ attains dimension-independent convergence if
@@ 6 Numerical experiments
This section tests our theoretical results on integrands with varying dimensions and decay rates of $\gamma_{u}$. We also reference [pan2024automatic], which provides further experiments demonstrating near-optimal convergence rates of median RQMC and comparisons to higher-order scrambled digital nets.
The test integrands have the common form
@@
For each integrand $f_{s,\gamma,\alpha}$, we evaluate the performance of the median estimator $\hat{\mu}^{(r)}_{E}$ for $m=1,2,\dots,16$ under two randomization schemes: the completely random design (CRD) and the random linear scrambling (RLS). The estimator $\hat{\mu}^{(r)}_{E}$ is computed as the sample median of 2r12r-1 independent realizations of $\hat{\mu}_{E}$, where $\hat{\mu}_{E}$ is defined as in equation (2) with precision $E=64$. The observed integration errors in our tested range of $m$ is much larger than $2^{-E}$, ensuring that $\hat{\mu}^{(r)}_{E}$ closely approximates $\hat{\mu}^{(r)}_{\infty}$. Consequently, the error bounds in Theorem 2, 3 and 4 apply to $\hat{\mu}^{(r)}_{E}$ with probability at least $1-2^{-r}$. We set $r=10$, limiting the failure probability per data point to below $0.1\%$. As a benchmark, we compute the standard RQMC estimator (STD), defined as the average of the 2r12r-1 $\hat{\mu}_{E}$ realizations used by RLS. The generating matrices $\mathcal{C}_{j}$ used by RLS and STD come from [joe:kuo:2008].
@@
Finally, we compare the integration errors under RLS and STD. Across the vast majority of test cases, RLS demonstrates superior performance over STD, reinforcing that the median RQMC approach (via RLS) is a more robust alternative to traditional RQMC (via STD). We also note that STD exhibits convergence rates comparable to those of RLS and substantially surpasses the theoretical $O(2^{-(3/2)m})$ rate derived from the root mean squared error when $\gamma=3$ or $4$. This discrepancy arises because each estimator $\hat{\mu}_{\infty}$ in STD is a special case of $\hat{\mu}^{(r)}_{\infty}$ with $r=1$, which satisfies the rapidly converging probabilistic error bounds in Theorems 3 and 4. However, averaging multiple replicates in STD compromises these bounds, as even a single outlier among the replicates can dominate the error and inflate its magnitude. This is the key mechanism that motivates our shift toward median RQMC approaches.
@@ 7 Discussion
A natural extension of our analysis is to integrands with discontinuities in their derivatives or function values. As discussed in Subsection 2.4, the requirement $f\in C^{(\alpha,\dots,\alpha)}([0,1]^{s})$ can be relaxed provided that the weak derivatives satisfy equation (5). Notably, the fractional Vitali variation framework inherently accommodates discontinuities. By developing a theory for non-smooth integrands, our results could be extended to real-world applications in finance and related fields. For instance, [griebel2017anova] showcases an infinite-dimensional integration problem where all ANOVA components are smooth despite discontinuities in the integrand’s derivatives, suggesting that non-smoothness may not fundamentally compromise convergence rates in high-dimensional settings.
Appendix A: Proof of Lemma 4
$|\hat{f}_{u}(\boldsymbol{k}^{\prime})|^{2}=$
where
In the special case v=1:sv=1{:}s, [pan2024automatic, page 14 line 14] shows
$\sum_{\kappa^{-}_{1},\dots,\kappa^{-}_{s}}\Bigl|\int_{[0,1]^{s}}g(\boldsymbol{x})\prod_{j\in 1{:}s}{\mathrm{wal}_{\kappa^{-}_{j}}}(x_{j})(-1)^{\vec{x}_{j}(\ell_{j})}W_{\kappa^{+}_{j}}(x_{j})\,\mathrm{d}\boldsymbol{x}\Bigr|^{2}\\ \leqslant 2^{-s}\Bigl(V^{(s)}_{\lambda}(g)\Bigr)^{2}\Bigl(\prod_{j=1}^{s}4^{-\lambda\ell_{j}}\prod_{\ell^{\prime}\in\kappa^{+}_{j}}4^{-\ell^{\prime}-1}\Bigr),$
where the summation $\sum_{\kappa^{-}_{1},\dots,\kappa^{-}_{s}}$ means each $\kappa^{-}_{j}$ runs over all subsets of $\mathbb{N}$ satisfying $\lceil\kappa^{-}_{j}\rceil_{1}<\ell_{j}$. After mapping $v$ to 1:|v|1{:}|v| and compactly rewriting the summation as $\sum_{\boldsymbol{k}^{\prime}\in B(\boldsymbol{k}_{u},v)}$, we obtain
and $\lceil\kappa_{j}\rceil_{\alpha+1}=0$ for $j\in u\setminus v$.
Appendix B: Proof of Lemma 5
we can apply Lemma 4 and equation (15) to get
we have for any $\theta\in(0,1)$ and $\rho_{\theta}=4^{-1+\theta}$
$g(t)=\Big(\sum_{N=1}^{\infty}t^{N}\Big)^{\alpha^{\prime}},$
and let $t^{N}]g(t)$ denote the coefficient of $t^{N}$ in $g(t)$ (see [flaj:sedg:2009] for background knowledge on generating functions). It follows that $t^{N}]g(t)$ gives the number of ways to choose an ordered array of length $\alpha^{\prime}$ from $\mathbb{N}$ while ensuring the elements in the array sum to $N$. Because each $\kappa\in\mathcal{S}_{\alpha^{\prime}}(\mathbb{N})$ is repeated $(\alpha^{\prime})!$ times in the above counting, we have
Our conclusion follows once we put the above bounds into equation (35).
Appendix C: Proof of Lemma 6
Next, we notice
$\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}2^{-|v|}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v,\\ \alpha_{j}>0\ \forall j\in u\setminus v\end{subarray}}\sum_{N=0}^{\lfloor T\rfloor}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|\sum_{\begin{subarray}{c}\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}\\ \|\boldsymbol{\ell}_{v}\|_{1}\leqslant(T-N)/(\alpha+\lambda)\end{subarray}}\hskip-10.00002pt2^{\|\boldsymbol{\ell}_{v}\|_{1}}.$
Let $\overline{N}=\lfloor(T-N)/(\alpha+\lambda)\rfloor$. For nonempty $v\subseteq u$ and integer $n\geqslant|v|$, there are $n-1\choose|v|-1$} number of $\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}$ satisfying $\|\boldsymbol{\ell}_{v}\|_{1}=n$. Hence, when $\overline{N}\geqslant|v|$,
$\sum_{\begin{subarray}{c}\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}\\ \|\boldsymbol{\ell}_{v}\|_{1}\leqslant(T-N)/(\alpha+\lambda)\end{subarray}}\hskip-10.00002pt2^{\|\boldsymbol{\ell}_{v}\|_{1}}=$ $\sum_{n=|v|}^{\overline{N}}{n-1\choose|v|-1}2^{n}\leqslant\sum_{n=|v|}^{\overline{N}}{\overline{N}-1\choose|v|-1}2^{n}\leqslant{\overline{N}-1\choose|v|-1}2^{\overline{N}+1}.$ (39)
When $\overline{N}<|v|$, no $\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}$ satisfies $\|\boldsymbol{\ell}_{v}\|_{1}\leqslant(T-N)/(\alpha+\lambda)$ and the above sum is trivially 0. Putting the above bound into equation (38), we get
To bound the sum over $N$, consider the generating function
$g(t)=\prod_{j\in u}\Big(\sum_{N=1}^{\infty}t^{N}\Big)^{\alpha_{j}}.$
for $t=2^{-1/(\alpha+\lambda)}$ and
$=$ $A_{\alpha,\lambda}^{|v|}B_{\alpha,\lambda}^{|u|-|v|},$
$\leqslant$ $2^{T/(\alpha+\lambda)}B^{|u|}_{\alpha,\lambda}$.
Putting the above bounds into equation (40), we finally get
[…]
$=$ $2^{T/(\alpha+\lambda)}\Big(B_{\alpha,\lambda}+\max\big(T/(\alpha+\lambda),1\big)A_{\alpha,\lambda}\Big)^{|u|}.$
Appendix D: Proof of Lemma 7
We adopt the notation used in Appendix Appendix C: Proof of Lemma 6. For $\boldsymbol{k}_{u}\in\mathcal{K}_{u,v}(T)$, we again consider the one-to-one correspondence between $\kappa^{+}_{j}$ and $\kappa^{+}_{j}-\ell_{j}$ for $j\in v$. Because $d\in\mathbb{N}$ and $d<\alpha+\lambda$, we must have $d\leqslant\alpha$. By $|\kappa^{+}_{j}|=\alpha\geqslant d$ for $j\in v$,
and
Let
Similar to equation (38), we can write
$|K_{u}(T)\cap K^{\prime}_{u}(T^{\prime})|$ (43)
$N_{*}=\frac{(\alpha+\lambda)T^{\prime}-dT}{\alpha+\lambda-d}.$
Therefore,
$\leqslant$ $\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0}\\ N_{*}<N\leqslant T\end{subarray}}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|{\lfloor(T-N)/(\alpha+\lambda)\rfloor-1\choose|v|-1}2^{\lfloor(T-N)/(\alpha+\lambda)\rfloor+1},$
Next, we notice
[…]
$=$ $D_{\alpha,\lambda,d,\theta^{\prime}}^{|u|},$
where $D_{\alpha,\lambda,d,\theta^{\prime}}$ is defined in equation (26).
@@ Acknowledgments
The author thanks David Krieg and Art Owen for valuable discussions and two anonymous reviewers for many helpful comments.